a cube with edge length of 10cm is inscribed in a sphere, what is the surface area of the sphere




1
Expert's answer
2022-05-09T07:10:02-0400

The diameter "d" of the sphere is


"d=\sqrt{(10)^2+(10)^2+(10)^2}=10\sqrt{3} (cm)"

The surface area "A" of the sphere is


"A=\pi d^2=\pi(10\sqrt{3})^2=300\pi({cm}^2)"

The surface area of the sphere is "300\pi\ {cm}^2."